Required information [The following information applies to the questions displayed below.] Gabi Gram started The Gram Co., a new business that began operations on May 1. The Gram Co. completed the following transactions during its first month of operations. May 1 G. Gram invested $44,500 cash in the company in exchange for its common stock. 1 The company rented a furnished office and paid $2,200 cash for May’s rent. 3 The company purchased $4,590 of office equipment on credit. 5 The company paid $730 cash for this month’s cleaning services. 8 The company provided consulting services for a client and immediately collected $5,200 cash. 12 The company provided $2,500 of consulting services for a client on credit. 15 The company paid $790 cash for an assistant’s salary for the first half of this month. 20 The company received $2,500 cash payment for the services provided on May 12. 22 The company provided $3,600 of consulting services on credit. 25 The company received $3,600 cash payment for the services provided on May 22. 26 The company paid $4,590 cash for the office equipment purchased on May 3. 27 The company purchased $80 of advertising in this month’s (May) local paper on credit; cash payment is due June 1. 28 The company paid $790 cash for an assistant’s salary for the second half of this month. 30 The company paid $300 cash for this month’s telephone bill. 30 The company paid $300 cash for this month’s utilities. 31 The company paid $1,600 cash in dividends to the owner (sole shareholder). Required: 1. Enter the amount of each transaction on individual items of the accounting equation. Do not determine new account balances after each transaction. (Enter the transactions in the given order. Enter reductions to account balances with a minus sign.)
Answer #1
Date Assets = Liabilities + Equity
Cash + Accounts Receivable + Office
Equipment
= Accounts Payable + Common Stock - Dividends + Revenue - Expenses
May 1 44500 + + = + 44500 - + -
1 -2200 + + = + - + - 2200
3 + + 4590 = 4590 + - + -
5 -730 + + = + - + - 730
8 5200 + + = + - + 5200 -
12 + 2500 + = + - + 2500 -
15 -790 + + = + - + - 790
20 2500 + -2500 + = + - + -
22 + 3600 + = + - + 3600 -
25 3600 + -3600 + = + - + -
26 -4590 + + = -4590 + - + -
27 + + = 80 + - + - 80
28 -790 + + = + - + - 790
30 -300 + + = + - + - 300
30 -300 + + = + - + - 300
31 -1600 + + = + - 1600 + -
Balance 44500 + 0 + 4590 = 80 + 44500 - 1600 + 11300 - 5190